# Confidential, Attestable, and Efficient Inter‑CVM Communication with Arm CCA

### v2

@@ 1 Introduction
Enabling CSM may appear straightforward, yet it is a challenging undertaking. Depending on the architecture, it may be infeasible without hardware modifications, for example, in AMD SEV-SNP (Section 8.2). Even systems that enable similar memory sharing between enclaves (user-space TEEs) fail to provide flexible and scalable mechanisms that fit the needs of modern ML systems. For instance, Plug-In Enclaves [li2021confidential] enable read-only shared enclaves, which can be used to effectively share static resources such as libraries in serverless applications. Cerberus [lee2022cerberus] introduces formal techniques to verify such a sharing with the same read-only access model. However, both designs are insufficient for applications that require extensive inter-process communication (IPC), where writes must be visible to the other side. Elasticlave [yu2022elasticlave], on the other hand, supports writable shared memory using RISC-V Physical Memory Protection (PMP). However, the underlying isolation mechanism inherently limits the number of shareable regions: the total number of memory partitions (protected or shared) (e.g., to 16) due to limitations in the underlying isolation technology (risc-v’s Physical Memory Protection), thus, restricting both the number of coexisting enclaves and the number of shared regions.
@@ 2 Background & Motivation
In this section, we review the foundational concepts required for the remainder of the paper. We begin with an overview of Arm CCA (Section 2.1), followed by a discussion of two key components of the architecture: the Realm Management Monitor (Section 2.2) and the attestation framework (Section 2.3). We conclude this section with the motivation behind this work (Section 2.4).
@@ 2.1 Arm CCA
Arm CCA extends the Armv9-A architecture with realm world [^1] In some references, the term execution environment is used instead of world; however, in this work, the two terms are used interchangeably. and root world, orthogonal to the existing normal world (NW) and secure world. Each 4 KB frame of physical memory (also referred to as a granule) is tagged with the world it belongs to at any given time. This ownership information is recorded in a structure called Granule Protection Table (GPT). A hardware mechanism called Granule Protection Check (GPC) enforces access restrictions based on both the ownership state of each granule (as recorded in the GPT) and the current processor state. A memory access is permitted only if it complies with the rules defined in Table I. In particular, when the processor operates in the root world state, it has access to the physical address space (PAS) of all other worlds. When operating in the realm or secure world, the processor can access the normal world’s PAS, but the realm and secure worlds are isolated from each other. The normal world has no access to the PAS of any other world. Arm CCA also leverages the isolation primitives of the Arm architecture—such as exception levels (ELs) and virtualization. Exception levels begin with EL3, the highest privilege level in the system, while EL2–EL0 provide intra-world privilege separation. The architecture support for virtualization includes two stages of address translation: mapping virtual addresses (VAs) to intermediate physical addresses [^2] We adopt Arm’s term intermediate physical address (IPA), which is also commonly referred to as guest physical address (GPA). (IPAs) and mapping IPAs to physical addresses (PAs).
@@ 3 System & Threat Model
In this section, we present system model (Section 3.1) and threat model (Section 3.2) of CAEC.
@@ 4 CAEC
In this section, we introduce CAEC. We begin by describing the setup of each realm in Section 4.1. We then present the full lifecycle of CSM in Section 4.2, and finally detail several key design components, including the realm identifier (Section 4.3) and physical memory management (Section 4.4).
@@ 4.1 Realm Setup
At a high level, CAEC extends Arm CCA with support for CSM, enabling realms to securely create, manage, and attach to CSM regions. The RMM serves as the trusted core of CAEC: it implements validation, enforces access control, and orchestrates all CSM-related operations. CAEC introduces new RSI commands to the RMM, which exposes the CSM-related services to realms (see Table II).
@@
Allocation Semantics. CAEC preserves the hypervisor’s control over physical memory management for both CSM and non-CSM regions, aligning with Arm CCA’s design principles. Since a CSM region is shared across multiple realms, memory-management flows must diverge from those used for regular private memory of realms. Accordingly, the RMM notifies the hypervisor whenever CSM regions are created, shared, or removed (through vCPU exits, see Table II). The hypervisor uses this information to populate the P-realm’s CSM region while ensuring that C-realms do not receive additional physical granules beyond what the P-realm explicitly shares (see Section 4.4 for further details). CAEC, however, guarantees that once a CSM region is established, all participating realms observe an identical and consistent view of the underlying physical granules.
@@
Realm Initialization. CAEC adapts the realm initialization flow of CCA, in which the hypervisor issues RMI commands [rmmspec] to populate the initial realm’s image and delegate realm’s metadata granules. The metadata granule introduced by CAEC (i.e., APT) is similarly created and initialized through a new RMI command, RMI_APT_CREATE (see Table II). At this stage, the APT is allocated for each realm within the RMM’s internal structures but does not yet contain any configuration. Once initialization is complete, the hypervisor begins scheduling the vCPUs of each realm. At this point, realms can communicate with external entities, including their respective realm owners. The realms establish TLS channel and prove them self to their realm owner via attestation (step 1 in Fig. 4).
@@ 4.2 CSM Lifecycle
Creation. A realm can create arbitrary CSM regions in its private address space via RSI_CSM_CREATE, providing the region’s start address and size (see Table II). While CAEC does not restrict the number of CSM regions a realm may create, each region must (1) be granule-aligned and (2) not overlap with any existing CSM region in the realm’s private address space. If these conditions are satisfied, the RMM registers this region within the APT of P-realm, notifies the hypervisor about the new CSM region via vCPU exit REC_EXIT_P_REALM_CSM. The hypervisor then populate the entire CSM range within the P-realm’s private address space (see Section 4.4). For every new delegated granule, the RMM creates the corresponding entry in the P-realm’s RTTs. The RMM later returns a unique region identifier IDCSM to the realm (steps 2.1 to 2.7 in Fig. 4).
Pre-Sharing. Before sharing CSM, realms must be able to securely identify their peers. To enable this, CAEC adapts the attestation primitives of CCA and delegates peer realm attestation to the realm owners. Each realm owner attests the peer realm to verify that the expected software stack is running within it (step  3.1). In CAEC, the RMM assigns each realm a unique identifier that is used to refer to that realm during subsequent CSM-related operations. The RMM also reports these identifiers as a separate claim in the attestation token. As a result, realm identifiers can be reliably recovered by realm owners during peer realm attestation. After validating the attestation report (i.e., checking that realm’s claims match the expected values), realm owners distribute the validated peer realm’s identifier to their respective realms (step 3.2). From this point onward, both realms know each others identifiers. In Section 4.3, we provide a detailed explanation on realm identifier and the way CAEC builds it.
@@ 4.4 Physical Memory Allocation
For CSM creation, the hypervisor performs the following steps for each granule in the CSM range. It first issues RMI_RTT_READ_ENTRY to check whether the target IPA is already populated. If it is not populated, the hypervisor issues RMI_GRANULE_DELEGATE followed by RMI_DATA_CREATE_UNKNOWN to delegate a physical granule and create the corresponding mapping in the realm’s RTT, respectively. For CSM attachment, the hypervisor again invokes RMI_RTT_READ_ENTRY. If the IPA is already populated, it reclaims the granule using RMI_DATA_DESTROY and RMI_GRANULE_UNDELEGATE. The hypervisor also ensures that the RTTs of both the P-realm and the C-realm exist for the CSM range. If necessary, it creates them by issuing RMI_GRANULE_DELEGATE followed by RMI_RTT_CREATE. As outlined in the adversary model (Section 3.2), we assume that the hypervisor correctly allocates all required resources to ensure uninterrupted realm execution. However, failure to allocate these granules does not compromise the security guarantees of CAEC.
@@ 5 Security Analysis
In this section we provide an in-depth analysis of different attacks vectors introduced in threat model (Section 3.2), explaining how CAEC remains protected against these threats.
CSM Protection. As discussed in Section 2.1, the hardware-enforced GPC mechanism ensures that any access to realm-world memory from either the normal world or the secure world is strictly prohibited. Since the RMM enforces that all CSM regions reside the realm world, CSM regions are inherently protected from direct access attempts originating from normal world and secure world actors.
@@ 6 Implementation
Software Stack. We use unmodified Trusted Firmware-A [TF-A] (v2.11) as the Monitor. We adapt the Trusted Firmware reference implementation of the RMM [RMM] (v0.5.0), extending it to support the new and modified commands introduced by CAEC (Table II). On the host side, we base our hypervisor on linux-cca [linux-cca] (v5+v7) and the virtual machine manager on kvmtool-cca [kvmtool-cca] (v3/cca). We extend the Linux KVM (Kernel-based Virtual Machine) module [kvm, dall2014kvm] to support the new RMI commands and to delegate physical memory to CSM regions in response to the newly defined vCPU exits from the RMM (Table II). We modify kvmtool to insert a node into the guest device tree, reserving a portion of its private address space to be used as the CSM. On the guest side, we add CSM driver to linux-cca [linux-cca] (v5+v7). This driver discovers the reserved region from the device tree, issues the appropriate RSI commands to create or attach to CSM regions, and exposes CSM to user space as a character device. The same driver is used by both P-realms and C-realms.
Table III summarizes the code changes introduced by CAEC, measured in lines of code (LoC). CAEC adds 1,062 LoC to the RMM (29k LoC), resulting in approximately a 4% increase in its size. The RMM version used in our prototype (v0.5.0) targets CCA v1.0, which lacks planned features such as device assignment and planes (Section 8.1). These forthcoming features are expected to significantly increase the RMM’s size, making CAEC ’s relative contribution to the TCB even smaller over time. On the host side, CAEC contributes 247 LoC to kvmtool, of which 213 LoC implement the PCI device used in our experimental setup (Section 7.3), and 34 LoC correspond to core CAEC functionality. CAEC also adds 394 LoC to the Linux KVM module. On the guest side, the CSM driver adds 467 LoC.
@@
Performance Prototype. At the time of writing, no commercial hardware implements the Arm CCA extensions. For performance prototyping, we therefore adopt OPENCCA [bertschi2025opencca], the first open-source performance prototype of Arm CCA. Our performance evaluation can be easily reproduced as OPENCCA runs on cost-effective hardware (i.e., Radxa Rock 5B [ROCK5B]). OPENCCA constructs the realm world as a separate execution environment within the normal world. Because current off-the-shelf hardware lacks essential CCA extensions (e.g., GPC), it cannot provide hardware-enforced isolated realm world. Nevertheless, performance prototyping remains feasible by emulating the realm world in software. This is achieved by splitting the normal-world software stack into two execution domains: one running the standard normal-world environment, and the other hosting the RMM and realm components. A modified Monitor then manages boot and context switching between these domains, effectively emulating transitions between the normal world and the realm world.
@@ 7 Evaluation
We answer Q1 in Section 7.2 and Q2 in Section 7.3, Section 7.4, and Section 7.5.
@@ 7.2 Compatibility with CCA hardware
To evaluate the compatibility of CAEC with real CCA hardware, we repeated the data-sharing benchmark (Section 7.4) on FVP. We observed no runtime errors or stalls on FVP cores during CSM creation, inference execution, and CSM termination, demonstrating that CAEC can operate on Arm CCA–enabled hardware. In all experiments, we use the CSM driver to expose the CSM region to user-space processes.
@@ 8 Discussion
In this section, we discuss three topics: why CAEC remains compatible with (and orthogonal to) future extensions of Arm CCA (Section 8.1); how CSM can be enabled in other confidential computing architectures (Section 8.2); and potential future research directions for CAEC (Section 8.3).
@@ Acknowledgment
We thank Jon Crowcroft for his invaluable suggestions that helped improve the paper. The research in this paper was supported by the UKRI Open Plus Fellowship (EP/W005271/1 Securing the Next Billion Consumer Devices on the Edge) and an Amazon Research Award “Auditable Model Privacy using TEEs”.
The authors employed ChatGPT models (GPT-5.1, GPT-5, and GPT-4o) as auxiliary tools for editorial support, exploration of related research, and code debugging. All generated material was examined and verified by the authors. The authors take full responsibility for the accuracy, integrity, and originality of the paper and the released code.

### v3

@@ 1 Introduction
Enabling CSM may appear straightforward, yet it is a challenging undertaking. Depending on the architecture, it may be infeasible without hardware modifications, for example, in AMD SEV-SNP (Sec. 8.2). Even systems that enable similar memory sharing between enclaves (user-space TEEs) fail to provide flexible and scalable mechanisms that fit the needs of modern ML systems. For instance, Plug-In Enclaves [li2021confidential] enable read-only shared enclaves, which can be used to effectively share static resources such as libraries in serverless applications. Cerberus [lee2022cerberus] introduces formal techniques to verify such a sharing with the same read-only access model. However, both designs are insufficient for applications that require extensive inter-process communication (IPC), where writes must be visible to the other side. Elasticlave [yu2022elasticlave], on the other hand, supports writable shared memory using RISC-V Physical Memory Protection (PMP). However, the underlying isolation mechanism inherently limits the number of shareable regions: the total number of memory partitions (protected or shared) is limited to the number of PMP registers (e.g., to 16), thus, restricting both the number of coexisting enclaves and the number of shared regions.
@@
  • All code of CAEC will be open-sourced upon acceptance.
@@ 2 Background & Motivation
In this section, we review the foundational concepts required for the remainder of the paper. We begin with an overview of Arm CCA (Sec. 2.1), followed by a discussion of two key components of the architecture: the Realm Management Monitor (Sec. 2.2) and the attestation framework (Sec. 2.3). We conclude this section with the motivation behind this work (Sec. 2.4).
@@ 2.1 Arm CCA
Arm CCA extends the Armv9-A architecture with realm world [^1] In some references, the term execution environment is used instead of world; however, in this work, the two terms are used interchangeably. and root world, orthogonal to the existing normal world (NW) and secure world. Each 4 KB frame of physical memory (also referred to as a granule) is tagged with the world it belongs to at any given time. This ownership information is recorded in a structure called Granule Protection Table (GPT). A hardware mechanism called Granule Protection Check (GPC) enforces access restrictions based on both the ownership state of each granule (as recorded in the GPT) and the current processor state. A memory access is permitted only if it complies with the rules defined in Tab. I. In particular, when the processor operates in the root world state, it has access to the physical address space (PAS) of all other worlds. When operating in the realm or secure world, the processor can access the normal world’s PAS, but the realm and secure worlds are isolated from each other. The normal world has no access to the PAS of any other world. Arm CCA also leverages the isolation primitives of the Arm architecture—such as exception levels (ELs) and virtualization. Exception levels begin with EL3, the highest privilege level in the system, while EL2–EL0 provide intra-world privilege separation. The architecture support for virtualization includes two stages of address translation: mapping virtual addresses (VAs) to intermediate physical addresses [^2] We adopt Arm’s term intermediate physical address (IPA), which is also commonly referred to as guest physical address (GPA). (IPAs) and mapping IPAs to physical addresses (PAs).
@@ 3 System & Threat Model
In this section, we present the system model (Sec. 3.1) and the threat model (Sec. 3.2) of CAEC.
@@ 4 CAEC
In this section, we introduce CAEC. We begin by describing the setup of each realm in Sec. 4.1. We then present the full lifecycle of CSM in Sec. 4.2, and finally detail several key design components, including the realm identifier (Sec. 4.3) and physical memory management (Sec. 4.4).
@@ 4.1 Realm Setup
At a high level, CAEC extends Arm CCA with support for CSM, enabling realms to securely create, manage, and attach to CSM regions. The RMM serves as the trusted core of CAEC: it implements validation, enforces access control, and orchestrates all CSM-related operations. CAEC introduces new RSI commands to the RMM, which exposes the CSM-related services to realms (see Tab. II).
@@
Allocation Semantics. CAEC preserves the hypervisor’s control over physical memory management for both CSM and non-CSM regions, aligning with Arm CCA’s design principles. Since a CSM region is shared across multiple realms, memory-management flows must diverge from those used for regular private memory of realms. Accordingly, the RMM notifies the hypervisor whenever CSM regions are created, shared, or removed (through vCPU exits, see Tab. II). The hypervisor uses this information to populate the P-realm’s CSM region while ensuring that C-realms do not receive additional physical granules beyond what the P-realm explicitly shares (see Sec. 4.4 for further details). CAEC, however, guarantees that once a CSM region is established, all participating realms observe an identical and consistent view of the underlying physical granules.
@@
Realm Initialization. CAEC adapts the realm initialization flow of CCA, in which the hypervisor issues RMI commands [rmmspec] to populate the initial realm’s image and delegate realm’s metadata granules. The metadata granule introduced by CAEC (i.e., APT) is similarly created and initialized through a new RMI command, RMI_APT_CREATE (see Tab. II). At this stage, the APT is allocated for each realm within the RMM’s internal structures but does not yet contain any configuration. Once initialization is complete, the hypervisor begins scheduling the vCPUs of each realm. At this point, realms can communicate with external entities, including their respective realm owners. The realms establish TLS channel and prove them self to their realm owner via attestation (step 1 in Fig. 4).
@@ 4.2 CSM Lifecycle
Creation. A realm can create arbitrary CSM regions in its private address space via RSI_CSM_CREATE, providing the region’s start address and size (see Tab. II). While CAEC does not restrict the number of CSM regions a realm may create, each region must (1) be granule-aligned and (2) not overlap with any existing CSM region in the realm’s private address space. If these conditions are satisfied, the RMM registers this region within the APT of P-realm, notifies the hypervisor about the new CSM region via vCPU exit REC_EXIT_P_REALM_CSM. The hypervisor then populate the entire CSM range within the P-realm’s private address space (see Sec. 4.4). For every new delegated granule, the RMM creates the corresponding entry in the P-realm’s RTTs. The RMM later returns a unique region identifier IDCSM to the realm (steps 2.1 to 2.7 in Fig. 4).
Pre-Sharing. Before sharing CSM, realms must be able to securely identify their peers. To enable this, CAEC adapts the attestation primitives of CCA and delegates peer realm attestation to the realm owners. Each realm owner attests the peer realm to verify that the expected software stack is running within it (step  3.1). In CAEC, the RMM assigns each realm a unique identifier that is used to refer to that realm during subsequent CSM-related operations. The RMM also reports these identifiers as a separate claim in the attestation token. As a result, realm identifiers can be reliably recovered by realm owners during peer realm attestation. After validating the attestation report (i.e., checking that realm’s claims match the expected values), realm owners distribute the validated peer realm’s identifier to their respective realms (step 3.2). From this point onward, both realms know each others identifiers. In Sec. 4.3, we provide a detailed explanation on realm identifier and the way CAEC builds it.
@@ 4.4 Physical Memory Allocation
For CSM creation, the hypervisor performs the following steps for each granule in the CSM range. It first issues RMI_RTT_READ_ENTRY to check whether the target IPA is already populated. If it is not populated, the hypervisor issues RMI_GRANULE_DELEGATE followed by RMI_DATA_CREATE_UNKNOWN to delegate a physical granule and create the corresponding mapping in the realm’s RTT, respectively. For CSM attachment, the hypervisor again invokes RMI_RTT_READ_ENTRY. If the IPA is already populated, it reclaims the granule using RMI_DATA_DESTROY and RMI_GRANULE_UNDELEGATE. The hypervisor also ensures that the RTTs of both the P-realm and the C-realm exist for the CSM range. If necessary, it creates them by issuing RMI_GRANULE_DELEGATE followed by RMI_RTT_CREATE. As outlined in the adversary model (Sec. 3.2), we assume that the hypervisor correctly allocates all required resources to ensure uninterrupted realm execution. However, failure to allocate these granules does not compromise the security guarantees of CAEC.
@@ 5 Security Analysis
In this section we provide an in-depth analysis of different attacks vectors introduced in threat model (Sec. 3.2), explaining how CAEC remains protected against these threats.
CSM Protection. As discussed in Sec. 2.1, the hardware-enforced GPC mechanism ensures that any access to realm-world memory from either the normal world or the secure world is strictly prohibited. Since the RMM enforces that all CSM regions reside the realm world, CSM regions are inherently protected from direct access attempts originating from normal world and secure world actors.
@@ 6 Implementation
Software Stack. We use unmodified Trusted Firmware-A [TF-A] (v2.11) as the Monitor. We adapt the Trusted Firmware reference implementation of the RMM [RMM] (v0.5.0), extending it to support the new and modified commands introduced by CAEC (Tab. II). On the host side, we base our hypervisor on linux-cca [linux-cca] (v5+v7) and the virtual machine manager on kvmtool-cca [kvmtool-cca] (v3/cca). We extend the Linux KVM (Kernel-based Virtual Machine) module [kvm, dall2014kvm] to support the new RMI commands and to delegate physical memory to CSM regions in response to the newly defined vCPU exits from the RMM (Tab. II). We modify kvmtool to insert a node into the guest device tree, reserving a portion of its private address space to be used as the CSM. On the guest side, we add CSM driver to linux-cca [linux-cca] (v5+v7). This driver discovers the reserved region from the device tree, issues the appropriate RSI commands to create or attach to CSM regions, and exposes CSM to user space as a character device. The same driver is used by both P-realms and C-realms.
Table III summarizes the code changes introduced by CAEC, measured in lines of code (LoC). CAEC adds 1,062 LoC to the RMM (29k LoC), resulting in approximately a 4% increase in its size. The RMM version used in our prototype (v0.5.0) targets CCA v1.0, which lacks planned features such as device assignment and planes (Sec. 8.1). These forthcoming features are expected to significantly increase the RMM’s size, making CAEC ’s relative contribution to the TCB even smaller over time. On the host side, CAEC contributes 247 LoC to kvmtool, of which 213 LoC implement the PCI device used in our experimental setup (Sec. 7.3), and 34 LoC correspond to core CAEC functionality. CAEC also adds 394 LoC to the Linux KVM module. On the guest side, the CSM driver adds 467 LoC.
@@
Performance Prototype. At the time of writing, no commercial hardware implements the Arm CCA extensions. For performance prototyping, we therefore adopt OPENCCA [bertschi2025opencca], the first open-source performance prototype of Arm CCA. Our performance evaluation can be easily reproduced as OPENCCA runs on a cost-effective hardware (i.e., Radxa Rock 5B [ROCK5B]). OPENCCA constructs the realm world as a separate execution environment within the normal world. Because current off-the-shelf hardware lacks essential CCA extensions (e.g., GPC), it cannot provide hardware-enforced isolated realm world. Nevertheless, performance prototyping remains feasible by emulating the realm world in software. This is achieved by splitting the normal-world software stack into two execution domains: one running the standard normal-world environment, and the other hosting the RMM and realm components. A modified Monitor then manages boot and context switching between these domains, effectively emulating transitions between the normal world and the realm world.
@@ 7 Evaluation
We answer Q1 in Sec. 7.2 and Q2 in Sec. 7.3, Sec. 7.4, and Sec. 7.5.
@@ 7.2 Compatibility with CCA hardware
To evaluate the compatibility of CAEC with real CCA hardware, we repeated the data-sharing benchmark (Sec. 7.4) on FVP. We observed no runtime errors or stalls on FVP cores during CSM creation, inference execution, and CSM termination, demonstrating that CAEC can operate on Arm CCA–enabled hardware. In all experiments, we use the CSM driver to expose the CSM region to user-space processes.
@@ 8 Discussion
In this section, we discuss three topics: why CAEC remains compatible with (and orthogonal to) future extensions of Arm CCA (Sec. 8.1); how CSM can be enabled in other confidential computing architectures (Sec. 8.2); and potential future research directions for CAEC (Sec. 8.3).
@@ Acknowledgment
We thank Jon Crowcroft for his invaluable suggestions that helped improve the paper. We also thank the anonymous reviewers for their insightful comments and suggestions. The research was supported by the UKRI Open Plus Fellowship (EP/W005271/1, Securing the Next Billion Consumer Devices on the Edge), the Amazon Research Award “Auditable Model Privacy using TEEs”, and the AI Security Institute (AISI) Systemic Safety Grants Programme (UKRI833).
Professor Kotz was supported by a Royal Society Wolfson Visiting Fellowship and by a collaborative award from the U.S. National Science Foundation (NSF) SaTC Frontiers program under award number 1955805.
The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of any sponsor. Any mention of specific companies or products does not imply any endorsement by the authors, by their employers, or by their sponsors.
Finally note that the authors used ChatGPT models (GPT-5.1, GPT-5, and GPT-4o) as auxiliary tools for editorial support, exploration of related research, and code debugging. All generated material was examined and verified by the authors, who take full responsibility for the accuracy, integrity, and originality of the paper and the released code.
Appendix A Data Availability
All data and code required to reproduce our evaluations will be open-sourced at the time of acceptance.